# Spin-glass dynamics: experiment, theory and simulation

### v1

@@ IV.4 The statics-dynamics dictionary
We discuss here the important issue of the quantitative connection between the non-equilibrium evolution of a spin glass (dynamics) and its equilibrium behavior (statics), following Baity-Jesi et al. [36]. Accurate numerical computations of the non-equilibrium fluctuation-dissipation ratio for the EA model have been the basis for establishing a statics–dynamics dictionary. This is of great help for computing controlled extrapolations.
A framework has been emerging in the past few years for theoretical explanations of what can be, and is, observed under actual experimental conditions. One needs to reach a theoretical understanding of what can be measured under the “best” experimental conditions. For example, it was found in the seminal experiments of Joh et al. [159] that the largest spin-glass correlation length one could reach in the laboratory was of order $50~{}a_{0}$, with $a_{0}$ the mean distance between magnetic moments. This was in 1998. Today, because of the availability of single crystals, one can do better, reaching $234~{}a_{0}$ [299, recall Sec. II.5]. The two settings, numerical simulations and experiments, are plagued by different limitations. If they can be overcome, a joint understanding becomes possible.
@@
The basis for our discussion is the very slow growth of the coherence length in the low-temperature phase of spin glasses (already discussed in detail in Sec. IV.2). Another crucial ingredient for this analysis is the stochastic stability discussed in Sec. III.1 and Sec. IV.3 because it allows the establishment of a quantitative relationship between dynamics and statics, a major issue. As discussed in Sec. IV.3, one can use stochastic stability to derive the GFDR [122, 201].
@@
$x(C,L)\equiv\int_{0}^{C}\text{d}q\;P_{L}\left(q\right)\equiv-\frac{\text{d}\mathcal{S}\left(C,L\right)}{\text{d}C}\,$ (37)
(recall Sec. III.1 and the finite size $P_{L}(q)$ shown in Fig. 19), the dynamic linear response function measured in off-equilibrium experiments by switching on a field $\delta H$ at time $t_{\mathrm{w}}$
$\chi_{L}\left(t,t_{\mathrm{w}}\right)\equiv\lim_{\delta H\rightarrow 0}\;\frac{\delta m(t+t_{\mathrm{w}})}{\delta H}\;,$ (38)
@@
Numerical simulations of a $3D$ EA spin glass on a system of linear size $L$ and volume $V=L^{3}$ can produce data that follows the relevant experimental protocol. One cools the simulated system from high temperatures $T$ down to, say, $T=0.64\,T_{g}$ in zero magnetic field, $H=0$. The system then evolves during the waiting time $t_{\mathrm{w}}$, generating a growing coherence length $\xi(t,t_{\mathrm{w}};H\!=\!0)$. At $t_{\mathrm{w}}$, one switches on a magnetic field $H$ and evaluates for further times the linear response function, $\chi_{L}(t,t_{\mathrm{w}})$, together with the correlation function $C_{L}(t,t_{\mathrm{w}})$. We denote the infinite-volume limit of different observables by omitting the subscript $L$. For $T<T_{\mathrm{g}}$ the GFDR in Eq. (36) has the form
@@
Figure 28: The “naive” scaling does not work when the correlation function becomes too small. It also does not work for short times when close to $q_{\text{\tiny EA}}$. Reproduced from Fig. 2 of Baity-Jesi et al. [36].
Taking the limit of large $L$ in the l.h.s. of Eq. (39) means keeping $L\gg\xi_{\mathrm{micro}}$ for the whole time of the experiment (or the simulation). However, in a numerical simulation, the order of the limits in the r.h.s. of Eq. (39), must be reversed. We reach equilibrium (i.e., infinite time) by using some unphysical dynamics (recall Sec. II.7) in a relatively small system. The function $\mathcal{S}\left[q=C_{L}(t,t_{\mathrm{w}}),L\right]$ is computed numerically by integrating the finite size probability distribution for the overlap $P_{L}(q)$. In other words, one uses numerical simulations to mimic experiments by relating the non-equilibrium response of a very large system at finite times to the equilibrium overlap for a much smaller system. In particular, one may find the effective equilibrium size $L_{\text{eff}}(t,t_{\mathrm{w}})$ as the value of $L$ such that $T\chi(t,t_{\mathrm{w}})=\mathcal{S}\left(q=C(t,t_{\mathrm{w}}),L\right)$. [^6] Recall that, if we keep $t$ and $t_{\mathrm{w}}$ fixed, the limit of infinite size can be reached for both $C$ and $\chi$ without much pain in the off-equilibrium experiment (or simulation). Unfortunately, the limit of large $L$ for $\mathcal{S}(q,L)$ is a lot more difficult to obtain.
@@
This is the method illustrated in Fig. 25, where $k=3.7$ was shown to work to scale the $C_{2+2}$ correlation function. When studying the response function, however, this simple statics-dynamics dictionary does not work, as shown in Fig. 28, where it is represented by the dotted lines. Even allowing a different value of $k$ for each value of $t_{\mathrm{w}}$ (solid lines) does not work. Again, both for short and for long times, we do not get a good description of the data. Our scaling hypothesis was, somehow, too naive.
@@
The line of reasoning presented here is based on the idea that the experimental aging response functions contain information about the Parisi order parameter $P(q)$, and is motivated by the fact that the extrapolation to infinite waiting times is cumbersome and sometimes mysterious. The scaling hypotheses (41) and (42), based on assuming that two different length and time scales have to be incorporated in the SDD, allow an orderly, systematic and fully successful extrapolation. When $\xi(t+t_{\mathrm{w}})\gg\xi(t_{\mathrm{w}})$ this dictionary disagrees with the simple SDD base on a single length scale, and correctly reports that, under these conditions, $L_{\text{eff}}\sim\xi(t+t_{\mathrm{w}})$.
V Off-equilibrium dynamics: non-linear response at a single temperature
@@ V.1 Experiments on films and the lower critical dimension
The use of thin-film spin glasses has generated some remarkable dynamical properties. At first glance, it would seem improbable that a sufficient number of spins were contained in a thin film of thickness, say, 2 or 3 nm, for measurement with even the most sensitive of instruments. The situation changed with the introduction of thin-film multilayers comprised of a thin spin-glass layer separated by a non-magnetic layer sufficiently thick to eliminate any communication between the spin-glass layers. These were pioneered by Kenning et al. [175, 172], who used thin layers of CuMn separated by thick Cu layers. The use of the same metal for solvent and separator eliminated the problem of strains at the interfaces.
@@
In order to account for the thin film geometry, Guchhait et al. [137] described the coherent volume as “pancake-like,” with a perpendicular length scale $\xi^{\perp}(t,T)$ and a parallel length scale $\xi^{\parallel}(t,T)$. They defined an effective correlation length,
@@
where $T_{\text{q}}$ is the quenched temperature of Fig. 31 and $\nu$ is the Ising critical exponent of a pure $3D$ model.
@@
Figure 31: Temperature dependence of the dynamic susceptibility for a CuMn 13.5 at.multilayer at different observation times. Reproduced from Fig. 1 of Sandlund et al. [258].
@@
Simulations (see Sec. V.2), however, have found that the parallel correlation length continues to grow after $t_{\text{co}}$. To probe this time regime would require measurements over time scales much larger than heretofore attained in the laboratory.
@@
Throughout this review, recall the glossary in Sec. I.3, we have considered several length scales in parallel. On the one hand, from theory and simulations, we have the coherence length $\xi_{\mathrm{micro}}$ extracted from the $C_{4}$ correlation function as explained in Sec. IV.2. Such microscopic detail is not accessible to experiments, where instead a $\xi_{\mathrm{Zeeman}}$ based on the Zeeman energy is measured, as described in Secs. 13 and II.6. On the other hand, while simulations have reached longer times with the advent of special-purpose computers, they are still not quite at the experimental scale. This section will show how to relate experimental and numerical length scales and, hence, move towards the joint quantitative analysis of experiment and simulation that will be treated in the following sections.
@@
and then $\xi_{\text{macro}}(t_{\mathrm{w}})=[N_{f}(t_{\mathrm{w}})]^{1/(D-\theta/2)}$.
@@
where the finite-$\xi_{\mathrm{micro}}$ dependence of $z$ is controlled by an exponent $\omega$. At the critical point, $\omega$ should be the exponent of the first correction to scaling [see 27]. In the spin-glass phase, it would correspond to the replicon exponent, $\omega=\theta$.
@@
Eq. (51) can be extrapolated for typical experimental values in thin films to find $Z_{\text{c}}(T)=z(T,\xi_{\text{film}})T/T_{\mathrm{g}}$. These extrapolations are shown in Fig. 33. The agreement with the experimental value is very good. A modification of Eq. (49) based on activated dynamics [66, 53] has less success in reproducing the nearly constant $Z_{\text{c}}(T)$ seen in experiments [34].
@@ V.3 Direct comparison of CuMn experiments with Janus II simulations
The advent of dedicated supercomputers, such as Janus II, and of single crystals of CuMn [299] opened up the possibility of combining numerical and experimental results for the first time. Given this opportunity, the initial challenge was to define observables accessible to both theory and laboratory. On the one hand, one has direct access to microscopic configurations through simulations. On the other hand, experiments measure the system’s response to external magnetic field perturbations, either switching it on (ZFC) or off (TRM). The previous section showed how these two approaches could be reconciled, by relating the experimental correlation length $\xi_{\mathrm{Zeeman}}$ to the microscopic coherence length $\xi_{\mathrm{micro}}$ measured in simulations. The key was the attempt by Baity-Jesi et al. [33] to measure experimental response functions in simulations. That approach, however, relied on an approximate scaling law, valid only for some range of values of the external field and which broke down close to $T_{\mathrm{g}}$ (Fig. 33), which is precisely the most interesting experimental regime for glassy systems (because it allows the observation of larger coherence lengths).
@@
Here $\xi$ stands for $\xi_{\mathrm{micro}}(t_{\mathrm{w}},T)$ [which it is equivalent to $\xi_{\mathrm{Zeeman}}(t_{\mathrm{w}},T)$, recall Fig. 32], $\hat{S}$ is a constant, $D=3$ is the spatial dimension and $\theta$ stands for the replicon exponent $\theta(\tilde{x})$, where $\tilde{x}=\ell_{\text{J}}(T)/\xi(t_{\mathrm{w}},T)$ and $\ell_{\text{J}}(T)$ is the Josephson length.
@@
(ii) However, the observation that the Taylor expansion
@@
where ${\hat{S}}(C_{\text{peak}})$ is the function appearing in the fluctuation-dissipation relations [89, 204, 122, 36, see Secs. IV.3 and IV.4]. From now on, we shall use the shorthand $\hat{S}$ for ${\hat{S}}(C_{\text{peak}})$.
Figure 36: The non-linear part of the response-time data, $[\ln(t^{\mathrm{eff}}_{H}/t^{\mathrm{eff}}_{H\to 0^{+}})-a_{2}(T)H^{2}]\xi^{\theta(\tilde{x})/2}$, plotted against the scaling variable $[H^{2}\xi^{D-\theta(\tilde{x})/2}]^{2}$, see Eq. (55). On the right, numerical data; on the left the experimental one. See Tables 1-2 for parameters details. Figure taken from Zhai et al. [302].
@@
$T_{\mathrm{m}}$ $t_{\mathrm{w}}$ $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ $t_{\text{max}}$ $\theta(\tilde{x})$ $C_{\mathrm{peak}}(t_{\mathrm{w}})$
@@
Figure 37: Growth of the rescaled magnetization for different experimental protocols: $[M_{\mathrm{ZFC}}(t,t_{\mathrm{w}})+M_{\mathrm{TRM}}(t,t_{\mathrm{w}})]/H$, and $M_{\mathrm{FC}}(0,t_{\mathrm{w}}+t)/H$. The lighter colors refer to the FC case; the darker ones correspond to the quantity ZFC + TRM. The vertical dashed lines indicate the effective times, $t^{\text{eff}}_{H}$ , associated with each case. The violation of Eq. (1) is evident. Figure taken from Paga et al. [231].
@@
[…]

### v2

@@ IV.4 The statics-dynamics dictionary
We discuss here the important issue of the quantitative connection between the nonequilibrium evolution of a spin glass (dynamics) and its equilibrium behavior (statics), following Baity-Jesi et al. [38]. Accurate numerical computation of the nonequilibrium fluctuation-dissipation ratio for the EA model has been the basis for establishing a statics–dynamics dictionary. This is of great help for computing controlled extrapolations.
A framework has been emerging in the past few years for theoretical explanations of what can be, and is, observed under actual experimental conditions. One needs to reach a theoretical understanding of what can be measured under the “best” experimental conditions. For example, it was found in the seminal experiments of Joh et al. [188] that the largest spin-glass coherence length one could reach in the laboratory was of order $50~a_{0}$, with $a_{0}$ the mean distance between magnetic moments. This was in 1998. Today, because of the availability of single crystals, one can do better, reaching $234~a_{0}$ [374, recall Sec. II.5]. The two settings, numerical simulations and experiments, are plagued by different limitations. If they can be overcome, a joint understanding becomes possible.
@@
The basis for our discussion is the very slow growth of the coherence length in the low-temperature phase of spin glasses (already discussed in detail in Sec. IV.2). Another crucial ingredient for this analysis is the stochastic stability discussed in Sec. III.1 and Sec. IV.3, because it allows the establishment of a quantitative relationship between dynamics and statics, a major issue. As discussed in Sec. IV.3, one can use stochastic stability to derive the GFDR [143, 240].
@@
$x(C,L)\equiv\int_{0}^{C}\text{d}q\;P_{L}\left(q\right)\equiv-\frac{\text{d}\mathcal{S}\left(C,L\right)}{\text{d}C}\,$, (41)
recall Sec. III.1 and the finite-size $P_{L}(q)$ shown in Fig. 20; the dynamic linear response function measured in off-equilibrium experiments by switching on a field $\delta H$ at time $t_{\mathrm{w}}$
$\chi_{L}\left(t,t_{\mathrm{w}}\right)\equiv\lim_{\delta H\rightarrow 0}\;\frac{\delta m(t+t_{\mathrm{w}})}{\delta H}\;$; (42)
@@
Numerical simulations of a $3D$ EA spin glass on a system of linear size $L$ and volume $V=L^{3}$ can produce data that follows a relevant experimental protocol. One cools the simulated system from high temperatures $T$ down to, say, $T=0.64\,T_{g}$ in zero magnetic field, $H=0$. The system then evolves during the waiting time $t_{\mathrm{w}}$, generating a growing coherence length $\xi(t,t_{\mathrm{w}};H=0)$. At $t_{\mathrm{w}}$, one switches on the magnetic field $H$ and evaluates the linear response function, $\chi_{L}(t,t_{\mathrm{w}})$ for further times, together with the correlation function $C_{L}(t,t_{\mathrm{w}})$. We denote the infinite-volume limit of different observables by omitting the subscript $L$. For $T<T_{\mathrm{g}}$ the GFDR in Eq. (40) has the form
@@
Figure 29: The “naive” scaling of Eq. (44) does not work when the correlation function becomes too small. It also does not work for short times when close to $q_{\text{\tiny EA}}$. Reproduced from Fig. 2 of Baity-Jesi et al. [38].
Taking the limit of large $L$ in the l.h.s. of Eq. (43) means keeping $L\gg\xi_{\mathrm{micro}}$ for the whole time of the experiment (or the simulation). However, in a numerical simulation, the order of the limits in the r.h.s. of Eq. (43), must be reversed. We reach equilibrium (i.e., infinite time) by using some unphysical dynamics (recall Sec. II.8) in a relatively small system. The function $\mathcal{S}\left[q=C_{L}(t,t_{\mathrm{w}}),L\right]$ is computed numerically by integrating the finite size probability distribution for the overlap $P_{L}(q)$. In other words, one uses numerical simulations to mimic experiments by relating the nonequilibrium response of a very large system at finite times to the equilibrium overlap for a much smaller system. In particular, one may find the effective equilibrium size $L_{\text{eff}}(t,t_{\mathrm{w}})$ as the value of $L$ such that $T\chi(t,t_{\mathrm{w}})=\mathcal{S}\left[q=C(t,t_{\mathrm{w}}),L\right]$. [^8] Recall that, if we keep $t$ and $t_{\mathrm{w}}$ fixed, the limit of infinite size can be reached for both $C$ and $\chi$ without much pain in the off-equilibrium experiment (or simulation). Unfortunately, the limit of large $L$ for $\mathcal{S}(q,L)$ is considerably more difficult to obtain.
@@
This is the method illustrated in Fig. 26, where $k=3.7$ was shown to work to scale the $C_{2+2}$ correlation function. When studying the response function, however, this simple statics-dynamics dictionary does not work, as shown in Fig. 29, where it is represented by the dotted lines. Even allowing a different value of $k$ for each value of $t_{\mathrm{w}}$ (solid lines) does not work. Again, we do not get a good description of the data, either for short or for long times. Our scaling hypothesis was, somehow, too naive.
@@
The line of reasoning presented here is based on the idea that the experimental aging response function contains information about the Parisi order parameter $P(q)$, and is motivated by the fact that the extrapolation to infinite waiting times is cumbersome and sometimes mysterious. The scaling hypotheses (45) and (46), based on assuming that two different length and time scales have to be incorporated in the SDD, allow an orderly, systematic and fully successful extrapolation. When $\xi(t+t_{\mathrm{w}})\gg\xi(t_{\mathrm{w}})$ this dictionary disagrees with the simple SDD base on a single length scale, and correctly reports that, under these conditions, $L_{\text{eff}}\sim\xi(t+t_{\mathrm{w}})$.
V Off-equilibrium dynamics: nonlinear response at a single temperature
@@ V.1 Experiments on films and the lower critical dimension
The use of thin-film spin glasses has generated some remarkable dynamical properties. At first glance, it would seem improbable that a sufficient number of spins are contained in a thin film of thickness, say, 2 or 3 nm, for measurement with even the most sensitive of instruments. The situation changed with the introduction of thin-film multilayers comprised of a thin spin-glass layer separated by a nonmagnetic layer sufficiently thick to eliminate any communication between the spin-glass layers. These were pioneered by Kenning et al. [209, 206], who used thin layers of CuMn separated by thick Cu layers. The use of the same metal for solvent and separator eliminated the problem of strains at the interfaces.
@@
In order to account for the thin-film geometry, Guchhait et al. [161] described the coherent volume as “pancake-like,” with a perpendicular length scale $\xi^{\perp}(t,T)$ and a parallel length scale $\xi^{\parallel}(t,T)$. They defined an effective correlation length,
@@
where $T_{\text{q}}$ is the quenched temperature of Fig. 32 and $\nu$ is the correlation-length critical exponent for the 2D{2D} Ising Edwards-Anderson model.
@@
Figure 32: Temperature dependence of the dynamic susceptibility for a CuMn 13.5 at.% multilayer at different observation times. Reproduced from Fig. 1 of Sandlund et al. [323].
@@
Simulations (see Sec. V.2), however, have found that the parallel correlation length continues to grow after $t_{\text{co}}$. Measurements at times greater than $t_{\text{co}}$ are an opportunity for future experiments (see Sec. VII.2).
@@
Throughout this review, recall the glossary in Sec. I.3, we have considered several length scales in parallel. On the one hand, from theory and simulations, we have the coherence length $\xi_{\mathrm{micro}}$ extracted from the $C_{4}$ correlation function as explained in Sec. IV.2. Such microscopic detail is not accessible to experiments, where instead a $\xi_{\mathrm{Zeeman}}$ based on the Zeeman energy is measured, as described in Secs. II.3 and II.6. On the other hand, while simulations have reached longer times with the advent of special-purpose computers, they are still not quite at the experimental scale. This section will show how to relate experimental and numerical length scales and, hence, move toward the joint quantitative analysis of experiment and simulation that will be treated in the following sections.
@@
with $N_{f}(t_{\mathrm{w}})$ the number of unpaired moments arising from fluctuations. Then, $\xi_{\text{macro}}(t_{\mathrm{w}})=[N_{f}(t_{\mathrm{w}})]^{1/(D-\theta/2)}$. The linear dependence of $E_{Z}$ upon $H$ would be expected at very small magnetic fields, while the quadratic dependence is expected for values of $H$ used in normal laboratory experiments.
@@
where $z_{\infty}(T)$ is the value of $z(T)$ in the limit that $\xi_{\text{micro}}\rightarrow\infty$, and the finite $\xi_{\mathrm{micro}}$ dependence of $z$ is controlled by the exponent $\omega$. At the critical point, $\omega$ should be the exponent of the first correction to scaling [see 29]. In the spin-glass phase, it would correspond to the replicon exponent, $\omega=\theta$.
@@
Eq. (55) can be extrapolated for typical experimental values in thin films to find $Z_{\text{c}}(T)=z(T,\xi_{\text{film}})T/T_{\mathrm{g}}$. These extrapolations are shown in Fig. 34. The agreement with the experimental value is very good. A modification of Eq. (53) based on activated dynamics [71, 56] has less success in reproducing the nearly constant $Z_{\text{c}}(T)$ seen in single-crystal experiments [36].
@@ V.3 Direct comparison of CuMn experiments with Janus II simulations
The advent of dedicated supercomputers, such as Janus II, and of single crystals of CuMn [374] opened up the possibility of combining numerical and experimental results for the first time. Given this opportunity, the initial challenge was to define observables accessible to both theory and laboratory. On the one hand, one has direct access to microscopic configurations through simulations. On the other hand, experiments measure the system’s response to external magnetic field perturbations, either switching it on (ZFC) or off (TRM). The previous section showed how these two approaches can be reconciled, by relating the experimental correlation length $\xi_{\mathrm{Zeeman}}$ to the microscopic coherence length $\xi_{\mathrm{micro}}$ measured in simulations. The key was the attempt by Baity-Jesi et al. [35] to measure experimental response functions in simulations. That approach, however, relied on an approximate scaling law, valid only for some range of values of the external field and which broke down close to $T_{\mathrm{g}}$ (Fig. 34), which is precisely the most interesting experimental regime for glassy systems (because it allows the observation of larger coherence lengths).
@@
[…]
@@
(ii) However, the Taylor expansion
@@
where ${\hat{S}}(C_{\text{peak}})$ is the function appearing in the fluctuation-dissipation relations [102, 243, 143, 38, see Secs. IV.3 and IV.4]. From now on, we shall use the shorthand $\hat{S}$ for ${\hat{S}}(C_{\text{peak}})$.
Figure 37: The nonlinear part of the response-time data, $\ln(t^{\mathrm{eff}}_{H}/t^{\mathrm{eff}}_{H\to 0^{+}})-a_{2}(T)H^{2}]\xi^{\theta(\tilde{x})/2}$, plotted against the scaling variable $H^{2}\xi^{D-\theta(\tilde{x})/2}]^{2}$, see Eq. (59). On the right, numerical data; on the left the experimental one. See Tables 1-2 for parameter details. Figure taken from Zhai et al. [376].
@@
[…]
@@
Figure 38: Growth of the rescaled magnetization for different experimental protocols: $M_{\mathrm{ZFC}}(t,t_{\mathrm{w}})+M_{\mathrm{TRM}}(t,t_{\mathrm{w}})]/H$, and $M_{\mathrm{FC}}(0,t_{\mathrm{w}}+t)/H$. Empty symbols are used for the FC case and filled symbols for the quantity ZFC + TRM. The vertical dashed lines indicate the effective times, $t^{\text{eff}}_{H}$ , associated with each case. The violation of Eq. (1) is evident. Figure taken from Paga et al. [288].
@@
$\xi_{\mathrm{micro}}(t_{\mathrm{w}})\propto H^{1/y_{\mathrm{H}}},\;\;2\,y_{\mathrm{H}}=D-\theta/2$ (64)